# Is Noise Conditioning Necessary for Denoising Generative Models?

### v1

@@ 6.2 Analysis
In Table 2, DDIM only fails with the deterministic ODE sampler (the default sampler in (Song et al., 2021a)); it still performs decently with the SDE sampler (i.e., the DDPM sampler). With this observation, we further investigate the level of stochasticity in Fig. 6.
Specifically, with the flexibility of DDIM (Song et al., 2021a), one can introduce a parameter $\lambda$ that interpolates between the ODE and SDE samplers by adjusting $\eta_{i}$ and $\zeta_{i}$ in Eq. 4 (see Eq. 63 in Section D.2). As shown in Fig. 6, increasing $\lambda$ (more stochasticity) consistently improves FID scores. When $\lambda=1$, DDIM behaves similarly to iDDPM.
@@
Thus far, our experiments have been on the CIFAR-10 class-unconditional task. To show the generalizability of our findings, we further evaluate class-unconditional generation on ImageNet $32{\times}32$, FFHQ $64{\times}64$, and class-conditional generation on CIFAR-10. See Table 3.
@@
Table 3: Changes of FID scores in the absence of noise conditioning, on class-unconditional ImageNet $32{\times}32$ and FFHQ $64{\times}64$, and class-conditional CIFAR-10.
  Model Sampler NFE FID w/ $t$rightarrow w/o $t$ ImageNet 32$\times$32   FM (1-RF) Euler 100 5.15 $\rightarrow$ 4.85 FFHQ 64$\times$64   EDM Heun 79 2.64 $\rightarrow$ 3.59 CIFAR-10 Class-conditional   EDM Heun 35 1.76 $\rightarrow$ 3.11 FM (1-RF) Euler 100 2.72 $\rightarrow$ 2.55
@@ 7 Discussion and Conclusion
We hope that rethinking the role of noise conditioning will open up new opportunities. Modern diffusion models are closely related to Score Matching (Hyvärinen & Dayan, 2005; Song & Ermon, 2019; Song et al., 2021b), which provides an effective solution to Energy-Based Models (EBM) (Hopfield, 1982; Ackley et al., 1985; LeCun et al., 2006; Song & Kingma, 2021). The key idea of EBM is to represent a probability distribution $p(x)$ by $p(x)=e^{-E(x)}/Z$, where $E(x)$ is the energy function. With the score function of $p(x)$ (that is, $\nabla_{x}E(x)$), one can sample from the underlying $p(x)$ by Langevin dynamics. This classical formulation models the data distribution $p(x)$ by a single energy function $E(x)$ that is solely dependent on $x$. Therefore, a classical EBM is inherently $t$-unconditional. However, with the presence of $t$-conditioning, the sampler becomes annealed Langevin dynamics (Song & Ermon, 2019), which implies a sequence of energy functions $\{E(x,t)\}_{t}$ indexed by $t$, with one sampling step performed on each energy. Our study suggests that it is possible to pursue a single energy function $E(x)$, aligning with the goal of classical EBM.
Our study also reveals that certain families of models, e.g., Flow Matching (Lipman et al., 2023; Liu et al., 2023; Albergo et al., 2023), can be more robust to the removal of $t$-conditioning. Although these models are closely related to diffusion, they can be formulated from a substantially different perspective—estimating a flow field between two distributions. While these models can inherit the $t$-conditioning design of diffusion models, their native formulation does not require the flow field to be dependent on $t$. Our study suggests that there exists a single flow field for these methods to work effectively.
@@ A.2 Numerical Experiments
Empirical (×$\times 10^{-4}$) Estimation (×$\times 10^{-4}$) Empirical Estimation Empirical
0.1 $0.0143\pm 0.0002$ $0.0163$ $0.558\pm 0.005$ $0.628$ $3894\pm 87$
@@ A.3 Evaluation of the bound values
In this section, we provide additional experiment details on how we compute the bound values and present the plot of the bound terms $A_{i}B_{i}$ in Fig. 4. For reference, we also include separate plots for $A_{i}$ and $B_{i}$ in Figs. 8(a) and 8(b)  [^6] An interesting fact in Fig. 8(b) is that, for EDM, there is a “phase change” at around $i=50$, which is caused by a non-smooth $\delta$ value. We hypothesize that this transition occurs at a noise level high enough that the data distribution can no longer be approximated as a point distribution, leading to a noticeable shift in behavior..
@@
(2) When the noisy image approaches clean data, some properties of the effective target $R$ become bad (e.g. the Lipchitz constant $L_{i}$ will be very big). This leads to a large error estimation, but in real cases, as the neural network will smooth the learned function, the error is typically smaller. If we consider ignoring the last 10 steps in our bound value, the bound will be in a reasonable range (approximately 10 for FM and uEDM, 140 for EDM, and $>10^{5}$ for DDIM).
@@ B.1 General Experiment Configurations
We implement our main code base using Google TPU and the JAX (Bradbury et al., 2018) platform, and run most of our experiments on TPU v2 and v3 cores. As the official codes are mostly provided as GPU code, we re-implemented most of the previous work in JAX. For the faithfulness of our re-implementation, please refer to Section B.3.
@@
For evaluation of the generative models, we calculate FID (Heusel et al., 2017) between 50,000 generated images and all available real images without any augmentation. We used the pre-trained Inception-v3 model provided by StyleGAN3 (Karras et al., 2021) and converted it into a model class compatible with JAX. As we have reproduced most of the results in Section B.3, we believe that our FID calculation is reliable.
@@
In the original work of ADM, Dhariwal & Nichol (2021) don’t provide result on the CIFAR-10 dataset in class-unconditional settings. In our implementation of ADM, we keep the main method of learning ${\bm{\epsilon}}$-prediction and the variance ${\bm{\Sigma}}$ simultaneously, but employ it on the class-unconditional CIFAR-10 task. Notice that this ADM formulation is also not included in our theoretical framework, but it still gives a reasonable result after removing noise conditioning (see Table 2).
@@
A table of selected important hyperparameters can be found in Table 5. For ICM and ECM we use the RAdam (Liu et al., 2020) optimizer, while for all other models we use the Adam (Kingma & Ba, 2015) optimizer. Also, we set the parameter $\beta_{2}$ to $0.95$ to stabilize the training process.
For all CIFAR-10 experiments, we used the architecture of NCSN++ in (Song et al., 2021b), with 56M parameters. For the ImageNet $32{\times}32$ experiment, we used the same architecture but a larger scale, with a total of 210M parameters.
@@
For the class-conditional CIFAR-10 experiments, we use exactly the same configurations and hyperparameters of EDM and FM with the unconditional generation case, except that we train the network with labels. For the conditioning on labels, we use the architecture in Karras et al. (2022). We do not apply any kind of guidance at inference time.
@@
For FFHQ $64{\times}64$ experiments, we directly use the code provided by Karras et al. (2022) and run it on 8 H100 GPUs. We keep all hyperparameters the same as the original code in the experiments. For the removal of noise-conditioning, we simply set the $c_{\text{noise}}$ variable in the code to zero.
@@
  Experiment Duration Warmup Epochs Batch Size Learning Rate EMA Schedule EMA Half-life Images Dropout iDDPM & ADM 100M 200 2048 $8{\times}10^{-4}$ EDM 50M 0.15 iDDPM(${\mathbf{x}}$-pred) 200M 200 2048 $1.2{\times}10^{-3}$ EDM 50M 0.15 DDIM 100M 200 512 $4{\times}10^{-4}$ EDM 50M 0.1 FM 100M 200 2048 $8{\times}10^{-4}$ EDM 50M 0.15 FM ImageNet32 256M 64 2048 $2{\times}10^{-4}$ EDM 50M 0 EDM 200M 200 512 $1{\times}10^{-3}$ EDM 0.5M 0.13 uEDM (ours) 200M 200 512 $4{\times}10^{-4}$ EDM 0.5M 0.2 ICM & ECM 400M 0 1024 $1{\times}10^{-4}$ Const(0.99993) - 0.3
@@ B.3 Our reimplementation faithfulness
Note that iDDPM (Nichol & Dhariwal, 2021) only reports results with 1000 and 4000 training steps. Due to computational constraints, we only evaluate the model with 500 sampling steps. Since our 500-step result has already been better than the 1000-step result reported in their work, we believe that we successfully reproduced the model.
@@
iDDPM (Nichol & Dhariwal, 2021) - 500 3.13 -
@@
EDM (Karras et al., 2022) Heun 35 1.99 1.97
1-RF (Liu et al., 2023) RK45 $\sim$127 2.53 2.58
iCT (Song & Dhariwal, 2024) - 2 2.59 2.46
@@
EDM (Karras et al., 2022) Heun 35 1.79 1.76
ImageNet 32$\times$32
FM (Lipman et al., 2023) Euler 100 5.15 -
FM (Lipman et al., 2023) RK45 $\sim$125 4.30 5.02
FFHQ 64$\times$64
EDM (Karras et al., 2022) Heun 79 2.64 2.39
@@ C.1 Proof of the Effective Target
Statement 1  (Concentration of $p(t|{\mathbf{z}})$).  Consider a single datapoint ${\mathbf{x}}\in[-1,1]^{d}$, ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}})$, $t\sim\mathcal{U}[0,1]$, and ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$ (as in Flow Matching). Given a noisy image ${\mathbf{z}}=(1-t_{*}){\mathbf{x}}+t_{*}{\bm{\epsilon}}$ produced by a given $t_{*}$, the variance of $t$ under the conditional distribution $p(t|{\mathbf{z}})$, is:
$\mathrm{Var}_{t\sim p(t|{\mathbf{z}})}[t]\approx\frac{t_{*}^{2}}{2d},$ (27)
under the situation that the data dimension $d$ satisfies $\frac{1}{d}\ll t_{*}$ and $\frac{1}{d}\ll 1-t_{*}$.
Derivation [^7] Here, “derivation” refers to an intuitive explanation rather than a formal proof. We will verify the validity of these statements empirically.. Consider a fixed ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$, where ${\mathbf{x}}$ is the fixed data point and ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}})$. Without loss of generality, we assume that ${\mathbf{x}}=(x_{1},x_{2},\ldots,x_{d})$ such that $-1\leq x_{d}\leq x_{d-1}\leq\ldots\leq x_{1}\leq 1$, and denote ${\bm{\epsilon}}=(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{d}),{\mathbf{z}}=(z_{1},z_{2},\ldots,z_{d})$.
To facilitate analysis, we introduce the following transformed variables [^8] For simplicity and clarity, we assume that $d$ is an even number, which is generally the case for images. However, the proof can be naturally extended to the case where $d$ is odd.:
$\begin{cases*}{\mathbf{x}}^{\prime}:=(x_{1}-x_{2},x_{3}-x_{4},\ldots,x_{d-1}-x_{d})\\ {\bm{\epsilon}}^{\prime}:=(\epsilon_{1}-\epsilon_{2},\epsilon_{3}-\epsilon_{4},\ldots,\epsilon_{d-1}-\epsilon_{d})\\ {\mathbf{z}}^{\prime}:=(z_{1}-z_{2},z_{3}-z_{4},\dots,z_{d-1}-z_{d})\end{cases*}$ (28)
and similarly define ${\bm{\epsilon}}_{*}^{\prime}$. Then, we have
${\mathbf{z}}^{\prime}=(1-t){\mathbf{x}}^{\prime}+t{\bm{\epsilon}}^{\prime}.$ (29)
Notice that ${\bm{\epsilon}}^{\prime}\sim\mathcal{N}(0,2I_{\frac{d}{2}})$ and
$\|{\mathbf{x}}^{\prime}\|^{2}=\sum_{i=1}^{\frac{d}{2}}(x_{2i-1}-x_{2i})^{2}\leq\sum_{i=1}^{\frac{d}{2}}2(x_{2i-1}-x_{2i})\leq\sum_{i=1}^{\frac{d}{2}}2(x_{2i-2}-x_{2i})=2(x_{0}-x_{d})\leq 4,$ (30)
where we defined $x_{0}:=1$ for convenience. As a result, we have
@@
$\|{\mathbf{z}}^{\prime}\|^{2}$ $=\left\|(1-t){\mathbf{x}}^{\prime}+t{\bm{\epsilon}}^{\prime}\right\|^{2}$
$=(1-t)^{2}\|{\mathbf{x}}^{\prime}\|^{2}+t^{2}\|{\bm{\epsilon}}^{\prime}\|^{2}+2t(1-t){\mathbf{x}}^{\prime}\cdot{\bm{\epsilon}}^{\prime}.$ (31)
Consider the three terms in Section C.2. By the concentration property of a Gaussian random variable in high dimension, we know that
$\|{\bm{\epsilon}}^{\prime}\|^{2}\approx 2\times\frac{d}{2}=d$ (32)
Thus, as $d\gg\frac{1}{t},\frac{1}{1-t}$, we know that the second term in Section C.2 is of the order $\Theta(d)$, while the first and third term have order $o(d)$. We can then conclude
$t\approx\frac{\|{\mathbf{z}}^{\prime}\|}{\|{\bm{\epsilon}}^{\prime}\|}.$ (33)
Notice that Eq. 33 can also be applied to $t_{*}$ and ${\bm{\epsilon}}_{*}$, leading to the approximation $\|{\mathbf{z}}^{\prime}\|\approx t_{*}\|{\bm{\epsilon}}_{*}^{\prime}\|$. As both $\|{\bm{\epsilon}}^{\prime}\|$ and $\|{\bm{\epsilon}}_{*}^{\prime}\|$ are highly concentrated around $\sqrt{d}$, it follows that $t$ concentrates at $t_{*}$.
Next, we aim to quantify the concentration of $\|{\bm{\epsilon}}^{\prime}\|$ to further bound the variance of $t$. Importantly, this property depends solely on the standard Gaussian distribution, which we encapsulate in the following lemma.

### v2

@@ 6.2 Analysis
In Table 2, DDIM only fails with the deterministic ODE sampler (the default sampler in (song2021ddim)); it still performs decently with the SDE sampler (i.e., the DDPM sampler). With this observation, we further investigate the level of stochasticity in Fig. 6.
Specifically, with the flexibility of DDIM (song2021ddim), one can introduce a parameter $\lambda$ that interpolates between the ODE and SDE samplers by adjusting $\eta_{i}$ and $\zeta_{i}$ in Eq. 4 (see Eq. 56 in Section D.2). As shown in Fig. 6, increasing $\lambda$ (more stochasticity) consistently improves FID scores. When $\lambda=1$, DDIM behaves similarly to iDDPM.
@@
Thus far, our experiments have been on the CIFAR-10 class-unconditional task. To show the generalizability of our findings, we further evaluate class-unconditional generation on ImageNet ${\times}32$, FFHQ ${\times}64$, and class-conditional generation on CIFAR-10. See Table 3.
@@
Table 3: Changes of FID scores in the absence of noise conditioning, on class-unconditional ImageNet ${\times}32$ and FFHQ ${\times}64$, and class-conditional CIFAR-10.
  Model Sampler NFE FID w/ $t$ $\rightarrow$ w/o $t$ ImageNet $\times32$   FM (1-RF) Euler 100 5.15 $\rightarrow$ 4.85 FFHQ $\times64$   EDM Heun 79 2.64 $\rightarrow$ 3.59 CIFAR-10 Class-conditional   EDM Heun 35 1.76 $\rightarrow$ 3.11 FM (1-RF) Euler 100 2.72 $\rightarrow$ 2.55
6.4 Classifier-Free Guidance
We further examine the impact of omitting noise conditioning when using classifier-free guidance (CFG) (dhariwal2021diffusion), a standard technique for significantly improving sample quality in class-conditional diffusion models.
Our experiments with CFG are conducted on the ImageNet $256{\times}256$ dataset on SiT (sit2024), which is a flow-matching variant of DiT (dit2023). For comparison, we train SiT-B/2 under the original paper’s configuration for both noise-conditional and unconditional model. At inference, we employ an Euler sampler with 250 steps and vary the CFG scale. See more details in Appendix B.
The results in Fig. 8 indicate that removing the noise conditioning incurs almost no degradation at different guidance scales, corroborating our analysis.
Figure 8: Classifier-free guidance results for SiT-B/2 on ImageNet $256{\times}256$. We use an Euler sampler with 250 steps. The best guidance scale in each setting is marked with a star. All other experimental details follow the original paper. Removing noise conditioning leads to almost no degradation regardless of guidance scale.
@@ 7 Discussion and Conclusion
We hope that rethinking the role of noise conditioning will open up new opportunities. Modern diffusion models are closely related to Score Matching (hyvarinen2005estimation; song2019ncsn; song2021scorebased), which provides an effective solution to Energy-Based Models (EBM) (hopfield1982neural; ackley1985learning; lecun2006tutorial; song2021train). The key idea of EBM is to represent a probability distribution $p(x)$ by $p(x)=e^{-E(x)}/Z$, where $E(x)$ is the energy function. With the score function of $p(x)$ (that is, $\nabla_{x}E(x)$), one can sample from the underlying $p(x)$ by Langevin dynamics. This classical formulation models the data distribution $p(x)$ by a single energy function $E(x)$ that is solely dependent on $x$. Therefore, a classical EBM is inherently $t$-unconditional. However, with the presence of $t$-conditioning, the sampler becomes annealed Langevin dynamics (song2019ncsn), which implies a sequence of energy functions $\{E(x,t)\}_{t}$ indexed by $t$, with one sampling step performed on each energy. Our study suggests that it is possible to pursue a single energy function $E(x)$, aligning with the goal of classical EBM.
Our study also reveals that certain families of models, e.g., Flow Matching (lipman2023flow; liu2023flow; albergo2023stochastic), can be more robust to the removal of $t$-conditioning. Although these models are closely related to diffusion, they can be formulated from a substantially different perspective—estimating a flow field between two distributions. While these models can inherit the $t$-conditioning design of diffusion models, their native formulation does not require the flow field to be dependent on $t$. Our study suggests that there exists a single flow field for these methods to work effectively.
@@
Acknowledgements
We greatly thank Google TPU Research Cloud (TRC) for granting us access to TPUs. Q. Sun, Z. Jiang, and H. Zhao are supported by the MIT Undergraduate Research Opportunities Program (UROP). We thank Cheng Jiang, Yiyang Lu, Mingyang Deng, Xingjian Bai, and Xianbang Wang for their comments and discussions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.2 Numerical Experiments
Empirical ($\times 10^{-4}$) Estimation ($\times 10^{-4}$) Empirical Estimation Empirical
0.1 $0143\pm 0.0002 0.0$1630.0163 $0.558\pm 0.005$ $0.628$ $3894\pm 87$
@@ A.3 Evaluation of the Bound Values
In this section, we provide additional experiment details on how we compute the bound values and present the plot of the bound terms $A_{i}B_{i}$ in Fig. 4. For reference, we also include separate plots for $A_{i}$ and $B_{i}$ in Figs. 9(a) and 9(b)  [^6] An interesting fact in Fig. 9(b) is that, for EDM, there is a “phase change” at around $i=50$, which is caused by a non-smooth $\delta$ value. We hypothesize that this transition occurs at a noise level high enough that the data distribution can no longer be approximated as a point distribution, leading to a noticeable shift in behavior..
@@
(2) When the noisy image approaches clean data, some properties of the effective target $R$ become bad (e.g. the Lipchitz constant $L_{i}$ will be very big). This leads to a large error estimation, but in real cases, as the neural network will smooth the learned function, the error is typically smaller. If we consider ignoring the last 10 steps in our bound value, the bound will be in a reasonable range (approximately 10 for FM and uEDM, 140 for EDM, and $10^{5}$ for DDIM).
@@ B.1 General Experiment Configurations
We implement our main code base using Google TPU and the JAX (jax2018github) platform, and run most of our experiments on TPU v2 and v3 cores. As the official codes are mostly provided as GPU code, we re-implemented most of the previous work in JAX. For the faithfulness of our re-implementation, please refer to Section B.3.
@@
For evaluation of the generative models, we calculate FID (heusel2017FID) between 50,000 generated images and all available real images without any augmentation. We used the pre-trained Inception-v3 model provided by StyleGAN3 (karras2021stylegan3) and converted it into a model class compatible with JAX. As we have reproduced most of the results in Section B.3, we believe that our FID calculation is reliable.
@@
In the original work of ADM, dhariwal2021diffusion don’t provide result on the CIFAR-10 dataset in class-unconditional settings. In our implementation of ADM, we keep the main method of learning $\bm{\epsilon}$}-prediction and the variance $\bm{\Sigma}$} simultaneously, but employ it on the class-unconditional CIFAR-10 task. Notice that this ADM formulation is also not included in our theoretical framework, but it still gives a reasonable result after removing noise conditioning (see Table 2).
@@
A table of selected important hyperparameters can be found in Table 5. For ICM and ECM we use the RAdam (Liu2020On) optimizer, while for all other models we use the Adam (kingma2014adam) optimizer. Also, we set the parameter $\beta_{2}$ to $0.95$ to stabilize the training process.
For all CIFAR-10 experiments, we used the architecture of NCSN++ in song2021scorebased, with 56M parameters. For the ImageNet ${\times}32$ experiment, we used the same architecture but a larger scale, with a total of 210M parameters. For experiments on classifier-free guidance on ImageNet ${\times}256$, we strictly follow the SiT-B/2 model with 400k training steps in sit2024.
@@
For the class-conditional CIFAR-10 experiments, we use exactly the same configurations and hyperparameters of EDM and FM with the unconditional generation case, except that we train the network with labels. For the conditioning on labels, we use the architecture in karras2022edm. We do not apply any kind of guidance at inference time.
@@
For FFHQ $64{\times}64$ experiments, we directly use the code provided by karras2022edm and run it on 8 H100 GPUs. We keep all hyperparameters the same as the original code in the experiments. For the removal of noise-conditioning, we simply set the $c_{\text{noise}}$ variable in the code to zero.
@@
  Experiment Duration Warmup Epochs Batch Size Learning Rate EMA Schedule EMA Half-life Images Dropout iDDPM & ADM 100M 200 2048 $8{\times}10^{-4}$ EDM 50M 0.15 iDDPM($\mathbf{x}$}-pred) 200M 200 2048 $1.2{\times}10^{-3}$ EDM 50M 0.15 DDIM 100M 200 512 $4{\times}10^{-4}$ EDM 50M 0.1 FM 100M 200 2048 $8{\times}10^{-4}$ EDM 50M 0.15 FM ImageNet32 256M 64 2048 $2{\times}10^{-4}$ EDM 50M 0 EDM 200M 200 512 $1{\times}10^{-3}$ EDM 0.5M 0.13 uEDM (ours) 200M 200 512 $4{\times}10^{-4}$ EDM 0.5M 0.2 ICM & ECM 400M 0 1024 $1{\times}10^{-4}$ Const(0.99993) - 0.3 SiT-B/2 100M 0 256 $1{\times}10^{-4}$ Const(0.9999) - 0
@@ B.3 Our Reimplementation Faithfulness
Note that iDDPM (nichol2021iddpm) only reports results with 1000 and 4000 training steps. Due to computational constraints, we only evaluate the model with 500 sampling steps. Since our 500-step result has already been better than the 1000-step result reported in their work, we believe that we successfully reproduced the model.
@@
iDDPM (nichol2021iddpm) - 500 3.13 -
@@
EDM (karras2022edm) Heun 35 1.99 1.97
1-RF (liu2023flow) RK45 $\sim$127 2.53 2.58
iCT (song2024improved) - 2 2.59 2.46
@@
EDM (karras2022edm) Heun 35 1.79 1.76
ImageNet $\times32$
FM (lipman2023flow) Euler 100 5.15 -
FM (lipman2023flow) RK45 $\sim$125 4.30 5.02
FFHQ $\times64$
EDM (karras2022edm) Heun 79 2.64 2.39
@@ C.1 Proof of the Effective Target
We claim in Fig. 3 that, for a fixed noisy image $\mathbf{z}$} whose true noise level is $t_{*}$, the posterior variance of $p(t|{\mathbf{z}})$ scales like $t_{*}^{2}/2d$. In this section, we first derive a $\mathcal{O}(t_{*}^{2}/d)$ upper bound on the variance under minimal technical assumptions. While obtaining the exact constant requires delicate optimizations, our Big-O presentation keeps the proof accessible. We then present a concise, intuitive argument to recover the $t_{*}^{2}/2d$ scaling, which— as confirmed by our numerical results in Section A.2—serves as an accurate and practical estimate of $\text{Var}_{p(t|{\mathbf{z}})}[t]$.
C.2.1 Rigorous Upper Bound on Variance.
Single Data Point Case.
We first consider the case where $N=1$. For brevity, write ${\mathbf{x}}={\mathbf{x}}_{1}$. Recall that our goal is to bound, with high probability over the randomness in $z\sim p(z)$, the posterior variance $\text{Var}_{t|{\mathbf{z}}}[t]$.
Note that some ill-behaved $\mathbf{z}$} can lead to strange distribution of $t$. Thus we work under the high-probability regime in which $\mathbf{z}$} concentrates around its typical behavior.
Because ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}}_{d})$ is rotationally invariant, we may, without loss of generality, assume $\mathbf{x}$} to only have $1$ nonzero coordinate.
Theorem 3.
Assume that
${\mathbf{x}}=(x,0,0,\cdots,0),x\in\left[-\sqrt{d},\sqrt{d}\right]$
${\mathbf{z}}=(1-t_{0}){\mathbf{x}}+t_{0}{\bm{\epsilon}},{\bm{\epsilon}}=(\epsilon_{0},{\bm{\epsilon}}^{\prime})$
Here, $\epsilon_{0}$ is a scalar that is the first coordinate of $\bm{\epsilon}$}, and $\bm{\epsilon}$$^{\prime}$ is its other coordinates.
We further assume that
@@
$t_{0}\in\left[\frac{1}{d},1-\frac{1}{d}\right],\quad||{\bm{\epsilon}}^{\prime 2}|-d|=\mathcal{O}(\sqrt{d}\log d),\quad|\epsilon_{0}|^{2}\leq\log d$ (27)
Then, we have $\mathrm{Var}_{t|{\mathbf{z}}}(t)=\mathcal{O}(t_{0}^{2}/d)$.
Note that Eq. 27 occur with probability $1-\mathcal{O}(\frac{1}{d})$, due to the norm-concentration properties of the Gaussian distribution.
Proof.
We directly compute the probability density functions as follows.
Firstly, we have
$p(t|{\mathbf{z}})\ \propto\ p({\mathbf{z}}|t)p(t)\ \propto\ \frac{1}{t^{d}}\exp\left(-\frac{1}{2}\left(\frac{|{\mathbf{z}}-(1-t){\mathbf{x}}|}{t}\right)^{2}\right)$
since the distribution $p({\mathbf{z}}|t)$ is simply a Gaussian with mean $(1-t){\mathbf{x}}$ and variance $t^{2}$ per dimension.
Looking into the exponent, it can be simplified as
$\left(\frac{|{\mathbf{z}}-(1-t){\mathbf{x}}|}{t}\right)^{2}=\left(\frac{|(t-t_{0}){\mathbf{x}}+t_{0}{\bm{\epsilon}}|}{t}\right)^{2}=\left|\left(1-\frac{t_{0}}{t}\right){\mathbf{x}}+\frac{t_{0}}{t}{\mathbf{z}}\right|^{2}$
$=$ $\left(\left(1-\frac{t_{0}}{t}\right)x+\frac{t_{0}}{t}\epsilon_{0}\right)^{2}+\left(\frac{t_{0}}{t}\right)^{2}|{\bm{\epsilon}}^{\prime}|^{2}$
which is a quadratic function on $\dfrac{t_{0}}{t}$. We write it as $a\left(\dfrac{t_{0}}{t}-\mu\right)^{2}+C$ for some constants $a,\mu,C$ independent of $t$, where
$a$ $=(x-\epsilon_{0})^{2}+|{\bm{\epsilon}}^{\prime}|^{2}=x^{2}-x\mathcal{O}(\sqrt{\log d})+d+\mathcal{O}(\sqrt{d}\log d)$
$=x^{2}+d+\mathcal{O}(\sqrt{d}\log d)$ (28)
and
$a\mu=x(x-\epsilon_{0})=x^{2}-\mathcal{O}(x\sqrt{\log d})$ (29)
by the assumption Eq. 27.
Substituting back, we get
$p(t|{\mathbf{z}})\ \propto\ \frac{1}{t^{d}}\exp\left(-\frac{1}{2}a\left(\frac{t_{0}}{t}-\mu\right)^{2}\right)$
We find its maximum by differentiating its log-density:
$\frac{\partial\log p\left(t|{\mathbf{z}}\right)}{\partial t}=-\frac{d}{t}+a\left(\frac{t_{0}}{t}-\mu\right)\frac{t_{0}}{t^{2}}=\frac{1}{t}\left(a\left(\frac{t_{0}}{t}\right)^{2}-a\mu\frac{t_{0}}{t}-d\right)$
Let $\lambda_{1}>0,\lambda_{2}<0$ be the two roots of the equation $f(X):=aX^{2}-a\mu X-d=0$ (since $a$ and $d$ are both positive).
Notice that $\lambda_{1}\lambda_{2}=-\dfrac{d}{a}$. We claim that
$\lambda_{1}=1+\mathcal{O}\left(\dfrac{\log d}{\sqrt{d}}\right),\lambda_{2}=-\dfrac{d}{a}\left(1+\mathcal{O}\left(\dfrac{\log d}{d}\right)\right)=-\Theta(1)$ (30)
This is because
$f(1)=\mathcal{O}(\sqrt{d}\log d),f^{\prime}(1)=x^{2}+2d+\mathcal{O}(\sqrt{d}\log d)$
by Section C.2.1 and Eq. 30, from which we derive the desired root bounds.
Using the factorization into roots, we have
$\frac{\partial\log p(t|{\mathbf{z}})}{\partial t}$ $=\frac{1}{t}\left(a\left(\frac{t_{0}}{t}\right)^{2}-a\mu\frac{t_{0}}{t}-d\right)=\frac{1}{t}a\left(\frac{t_{0}}{t}-\lambda_{1}\right)\left(\frac{t_{0}}{t}-\lambda_{2}\right)$
$=\frac{1}{t}\left(\frac{t_{0}}{t}-\lambda_{1}\right)\Omega(d)$
by Section C.2.1 and Eq. 30.
Denote $t^{*}=\dfrac{t_{0}}{\lambda_{1}}$ as the unique maximizer of $\log p(t|{\mathbf{z}})$. As a first step, we show that in range $(0,3t^{*}]$, the probability density of $p(t|{\mathbf{z}})$ is more concentrated than a Gaussian with variance $\mathcal{O}(t_{0}^{2}/d)$. Intuitively, it is due to the fact that in this range,
$\frac{\partial\log p(t|{\mathbf{z}})}{\partial t}=\frac{1}{t}\left(\frac{t_{0}}{t}-\frac{t_{0}}{t^{*}}\right)\Omega(d)=\frac{t_{0}(t^{*}-t)}{t^{2}t^{*}}\Omega(d)=(t^{*}-t)\Omega\left(\frac{d}{t_{0}^{2}}\right)$
utilizing Eq. 30. This inequality implies that in this range, the gradient of this probability density is sharper than $\mathcal{N}(t^{*},\mathcal{O}(t_{0}^{2}/d))$.
To be specific, when $0<t<t^{\prime}\leq t^{*}$ or $t^{*}\leq t^{\prime}<t\leq 3t^{*}$, we have
$\log p(t^{\prime}|{\mathbf{z}})-\log p(t|{\mathbf{z}})=\int\limits_{t}^{t^{\prime}}\frac{\partial\log p(t|{\mathbf{z}})}{\partial t}\mathrm{d}t=\Omega\left(\frac{d}{t_{0}^{2}}\right)\cdot\left((t-t^{*})^{2}-{(t^{\prime}-t^{*})}^{2}\right)$ (31)
where the RHS is exactly the difference in log-probability density of $\mathcal{N}(t^{*},\mathcal{O}(t_{0}^{2}/d))$ at $t$ and $t^{\prime}$. This fact supports that in this range, $p(t|{\mathbf{z}})$ is sharper than $\mathcal{N}(t^{*},\mathcal{O}(t_{0}^{2}/d))$, or
$\mathbb{E}_{t\sim p^{\prime}}[(t-t^{*})^{2}]\leq\mathbb{E}_{t\sim q}[(t-t^{*})^{2}]=\mathcal{O}(t_{0}^{2}/d)$
where $p^{\prime},q$ denotes the distribution of $p(t|{\mathbf{z}})$ and $\mathcal{N}(t^{*},\mathcal{O}(t_{0}^{2}/d))$ restricted on $(0,3t^{*}]$, respectively.
Finally, we only need to consider the part when $3t^{*}<t\leq 1$, where we are going to prove that
$\Pr_{t\sim p(t|{\mathbf{z}})}[t>3t^{*}]$
is small. In this case, according to Eq. 31, for any $t^{\prime}\in[t^{*},2t^{*}]$:
$\frac{p(t|{\mathbf{z}})}{p(t^{\prime}|{\mathbf{z}})}\leq\exp\left(-\Omega\left(\frac{d}{t_{0}^{2}}\right)\cdot(4{t^{*}}^{2}-{t^{*}}^{2})\right)=\exp(-\Omega(d))$
which is actually exponentially small. This implies that
$\Pr_{t\sim p(t|{\mathbf{z}})}[t>3t^{*}]\leq\exp(-\Omega(d))\cdot\min_{t^{\prime}\in[t^{*},2t^{*}]}p(t^{\prime}|{\mathbf{z}})\leq\exp(-\Omega(d))\cdot\frac{1}{t^{*}}=\mathcal{O}\left(\frac{1}{d^{3}}\right)$
Consequently, the contribution of this tail to the variance is $\mathcal{O}(1/d^{3})$, which is negligible compared to $\mathcal{O}(t_{0}^{2}/d)$. Combining the two regimes, we conclude that
$\text{Var}_{p(t|{\mathbf{z}})}[t]\leq\mathbb{E}_{t\sim p(t|{\mathbf{z}})}[(t-t^{*})^{2}]=\mathcal{O}(t_{0}^{2}/d)$
Multiple Data Points Case.
We now turn to the setting of $N>1$ data points. Intuitively, one can identify the ground-truth clean image that generates the given noisy image, as long as the noise is not too large to swamp all signal. This can be done simply by comparing inner product: the noisy observation $\mathbf{z}$} will correlate most strongly with its corresponding clean sample ${\mathbf{x}}_{i}$, and only weakly with all others.